OpenStudy (anonymous):

Solve the system of equations using the substitution method. Enter the answer as an ordered pair. x + y = 5 y = 7

OpenStudy (vijay):

-2,7

OpenStudy (anonymous):

x + (7) = 5 x = 5 - 7 x = -2 x,y = -2, 7

OpenStudy (anonymous):

y = 7 x = -2